OpenStudy (anonymous):

A taco stand sells tacos for $3.25 each. The stand’s expenses for the day are $210. Write an inequality to represent the amount of tacos they need to sell per day to make a profit, and then solve the inequality.

OpenStudy (anonymous):

.if they have to sell n tacos to make a profit, then 3.25*n>=210 ...

OpenStudy (anonymous):

im still confused....

OpenStudy (anonymous):

on what,?

OpenStudy (anonymous):

the asnsweerr...

OpenStudy (anonymous):

3.25x>210 (it needs to be greater than so they can make a profit) Divide from 3.25 from both sides: x>64.62, but if the problem asks you to round, then you could just say 65. You should display this on a number line.

OpenStudy (anonymous):

Figure out how many tacos they need to sell to make more than $210. n= expenses/cost of taco

OpenStudy (anonymous):

can you guys just give me the answer ill give you a medal

OpenStudy (anonymous):

i already gave u the answer!

OpenStudy (anonymous):

Okay... let's see.. In order to make at least $210 by the end of the day, they'd have to sell at least 65 tacos at $3.25 each (I divided $210 by $3.25 and rounded 64.6 up to 65 because you can't sell half of a taco). So, the inequality would be: \[n \ge 65\] n = number of tacos that must be sold.

OpenStudy (anonymous):

exactly what i said.^-__^

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

SO CAN SOME ONE JUSTWRITEE HOW TO WRITE AS AN INEQUALITY AND THE ANSWER...

OpenStudy (anonymous):

i already did it for u!

OpenStudy (anonymous):

I just wrote the inequality...

OpenStudy (anonymous):

look back @ my answer.

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

3.25x>210 (it needs to be greater than so they can make a profit) Divide from 3.25 from both sides: x>64.62, but if the problem asks you to round, then you could just say 65. You should display this on a number line.

OpenStudy (anonymous):

\[n \ge 65\] ^^^ there it is again.

OpenStudy (anonymous):

.if they have to sell n tacos to make a profit, then 3.25*n>=210 .

OpenStudy (anonymous):

and the answer....

OpenStudy (anonymous):

n>65

OpenStudy (anonymous):

lolllll oh thanks!

OpenStudy (anonymous):

make sure your inequality sign is "greater than or equal to", not just "greater than"

OpenStudy (anonymous):

yeah sorry @TammisaurusRex i meant that sign but i didn't know how to do it.

OpenStudy (anonymous):

i cant make the line on the sheet im doing....

OpenStudy (anonymous):

oh so then explain it in words.

OpenStudy (anonymous):

are you doing it on microsoft word?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Go to the top option bar, click "insert", and then click "symbol", and you should be able to find the "greater than or equal to" symbol.